If dividend payout ratio is increased to 100%, what happens to retained earnings?


A firm has a debt 10Mm They are loan ophm) Th first ortinn ís闷Syr 104n @ 8% fle+ ra due at the end of the Ss a lanu0 nce and tuis ottim ist yu io A 13 13/ Fowth yur Fifth y
Answer #1
As the compensating balance would also have to
be borrowed, to get S10M, the borrowing should
be for 10/90% = $         11.11 million
The effective rate of interest for the option will be:
= 11.11*6%/10 = 6.67%
This interest rate will be applicable for all the 5 years.
Hence, the second option is better.